# IMPROVE: Impact of Mobile Phones on Remote Online Virtual Education

### v1

@@ Abstract
This work presents the IMPROVE dataset, designed to evaluate the effects of mobile phone usage on learners during online education. The dataset not only assesses academic performance and subjective learner feedback but also captures biometric, behavioral, and physiological signals, providing a comprehensive analysis of the impact of mobile phone use on learning. Multimodal data were collected from 120 learners in three groups with different phone interaction levels. A setup involving 16 sensors was implemented to collect data that have proven to be effective indicators for understanding learner behavior and cognition, including electroencephalography waves, videos, eye tracker, etc. The dataset includes metadata from the processed videos like face bounding boxes, facial landmarks, and Euler angles for head pose estimation. In addition, learner performance data and self-reported forms are included. Phone usage events were labeled, covering both supervisor-triggered and uncontrolled events. A semi-manual re-labeling system, using head pose and eye tracker data, is proposed to improve labeling accuracy. Technical validation confirmed signal quality, with statistical analyses revealing biometric changes during phone use.
@@
Education is a fundamental societal pillar and has continually evolved throughout history. In recent years, it has been deeply influenced by the digital revolution, which has moved our society from the physical to the digital world. Evidence of this change in recent decades is the growing significance of e-learning or online education, expected to grow exponentially over the next 20 years [1, 2]. Relevant institutions are adopting this model, especially those offering MOOCs (Massive Open Online Courses), due to their ability to reach a broader learner base [3].
@@
Learners have also been significantly impacted by this digital revolution, altering their habits to incorporate electronic devices into their daily routines such as mobile phones and tablets. Since 2012, studies [4, 5] indicated that between 97%-99% of learners own a mobile phone. Moreover, 95% of learners bring their phones to class every day, 92% use these devices to send text messages during class, and 10% even admit to texting during exams. On average, learners spend about 5 hours per day on their mobile phones, predominantly using messaging services (sending an average of 109 text messages per day) or social networks like WhatsApp and Instagram [6, 7].
@@
Therefore, the mobile phone raises numerous questions: Can it become a distraction to learners’ attention? Or conversely, could the absence of a mobile phone cause greater distraction or anxiety? Considering that attention is defined as the ability to concentrate, specifically, to exert conscious cognitive effort on a particular task or stimulus at any given moment [8], and given that various studies suggest it is easy to lose focus due to short-term distractions such as messages, sounds, etc. [9, 10]; it is valid to consider how mobile phones could impact learner performance. This is not a new idea, and various studies have shown that mobile phone use negatively affects learners’ attention, memory, and academic performance. This is mainly because it increases cognitive load by inducing learners to multitask [10, 5, 11, 12]. This effect becomes more noticeable as the class progresses or in long-duration activities [11]. However, learners’ performance can be influenced by their emotional state. For this reason, nomophobia, defined as the fear, stress, or anxiety caused by being unable to access or use a mobile phone, is an important factor to consider. Various studies have demonstrated how this condition affects memory and attention, ultimately impacting learners’ academic performance [13, 11].
@@
It is interesting to note that, although nomophobia can lead to lower academic performance, the previously mentioned studies have shown that just the presence of a mobile phone also reduces learners’ attention levels, presenting a complex challenge that requires in-depth investigation.
@@
The IMPROVE database was obtained in a realistic online learning environment. A total of 120 learners were monitored, divided into three different groups to understand the effects of mobile phone usage. To our knowledge, IMPROVE dataset is the first to provide a comprehensive multimodal acquisition setup including behavioral [1, 14], biometric, and learning analytics data from learners during an authentic online course that considers mobile phone usage. The database aims to facilitate understanding of the effects of mobile phone usage in online education. Additionally, it provides a robust dataset that can be utilized to develop new technologies to enhance learners’ learning processes or ensure secure certification [14, 15, 16, 17, 18].
@@
We employed 16 sensors, recognized as effective indicators for understanding learner behavior [19, 8, 20, 21], including an electrodermal activity (EDA) sensor, an electroencephalogram (EEG) band, 2 pulse meters, an accelerometer, a gyroscope, 3 RGB cameras, 2 Near-Infrared (NIR) cameras, a mouse, a keyboard, a microphone, a monitor, and an eye tracker (see Fig. 1). These sensors enabled us to monitor biometric and behavioral signals throughout 120 online learning sessions, each lasting approximately 30 minutes. During these sessions, learners interacted with educational content related to HTML, including video lectures, reading documents, and completing tests to assess their knowledge both before and after the session (with a total of 16 questions per session). This comprehensive setup allowed us to track not only how learners engaged with the material but also how external factors, such as mobile phone interruptions, affected their focus and learning outcomes.
@@
In total, the database comprises 2.83 terabytes of data, making it one of the most extensive multimodal datasets in this domain. This includes 145 labeled mobile phone interruptions, providing crucial insights into how mobile phone usage affects learners.
@@
Figure 1: Acquisition setup during the data capture using the edBB[1, 14] platform, illustrating all the sensors and devices employed, along with examples of the captured information.
@@ Methods
This section outlines the experimental methods and materials utilized for data collection. The database was obtained in accordance with the Declaration of Helsinki and was approved by the Ethics Committee of the Universidad Autónoma de Madrid (No. 11/04/2023). Furthermore, the learners were informed about the experimental procedure and data acquisition setup, and they signed a written consent form. An informed consent form was used to inform the learners that the obtained information will be used solely for research purposes. Learners received an incentive for participating.
Database: IMPROVE
The database consists of 120 learners from the School of Engineering at the Universidad Autónoma de Madrid (UAM), who were monitored during a real online learning session focused on HTML lasting between 20 and 35 minutes. The session was part of the MOOC called “Introduction to Development of Web Applications” (WebApp MOOC for short) available on the edXaaahttps://www.edx.org/ platform.
@@
Table 1: Distribution of Learners by Degree Program, HTML Proficiency Levels, Medical Conditions (percentage are not shown because learners can have more than one condition), and Gender, with Average Ages.
@@
Learners from the School of Engineering at the Universidad Autónoma de Madrid (UAM) were recruited. The School of Engineering was selected because the five different degrees offered require basic computer science knowledge, which aligns well with the online course. To encourage participation, an economic incentive was provided. A total of 170 learners signed up for the study, from which 120 were selected. An initial test was conducted with 6 learners to verify the protocol and monitoring system. Adjustments were made after the initial test, such as: requiring Group 1 learners to provide more detailed responses to messages (for more information about the groups formed, see the subsection Protocol), providing a more detailed explanation of the session to learners before starting, eliminating an initial form to avoid potential suggestion, and making adjustments to the EEG band and the calibration of the eye tracker. The 6 learners from the initial test did not participate in the final data collection. 4 other learners were discarded due to issues during the session and replaced by other participants. The final 120 learners were divided into three groups of 40. The selection process considered various criteria to ensure balanced groups, including gender (61 male, 59 female), HTML proficiency (none, beginner, intermediate, advanced), degree program, among other factors. The learners’ age range was 18 to 32 years old, with an average age of 21 years old (see Table 1). The learners gave their consent for the publication of the database. Each learner was anonymized with an identifier that cannot be linked to their identity.
@@
The objective was to understand the effects of mobile phone use or absence on learner behavior and academic performance. To this end, three different groups were formed:
@@
    Group 1: Mobile phone use and possession allowed. The device was placed on the learner’s desk, visible to the learner, with sound and vibration activated.
@@
    Group 2: Mobile phone possession was allowed, but their use was prohibited. The device was also placed on the learner’s desk, but with the screen facing down, and with sound and vibration activated.
@@
    Group 3: The mobile phone was confiscated during the whole learning session.
Figure 2: Protocol during the learning session to evaluate the impact of mobile phone use on learners’ academic performance. The top section outlines the session protocol, dividing learners into three groups based on their allowed level of interaction with mobile phones. The bottom section illustrates the activities performed by learners and shows how events and activities were labeled for later analysis. A graph corresponding to a user’s attention signal is shown, with clear event labeling aligned with specific activities.
@@
Three supervisors were present throughout the session. The first supervisor was responsible for explaining the session in detail to the learner, describing the operation of the sensors, placing and calibrating them, and ensuring that the platforms were working correctly. The second supervisor used the LOGGE tool [22] (for more details on this tool, see the "Data, platforms and sensors" subsection) to label the moments when the learner interacted with the mobile phone or when it rang, as well as to label other relevant information for the session. The third supervisor was responsible for sending two messages to the learners during key moments of the learning session of both Group 1 and Group 2, reviewing the information recorded by the second supervisor, noting any possible incidents during the session, and assisting the learner in case of doubts. We will differentiate between two different types of interruptions: interventions (the two controlled messages sent by the supervisor) and casual interactions (uncontrolled messages/calls received by the learner out of the interventions). Therefore, interventions are considered controlled events, while interactions are classified as uncontrolled events. The first intervention was received during an explanatory video of the lesson, and the second during the reading of HTML code, which was part of the learner evaluation (see Fig. 2). Learners in Group 1 were required to respond with a detailed reply. The messages were simple questions such as: "What do you think of the cafeteria of our School of Engineering?" or an image of a popular movie like Harry Potter accompanied by the question: "What do you think of this film series? Would you recommend it?" On average, learners took about 26 seconds to respond to the first message and about 36 seconds to respond to the second message. Learners in Group 2 were not allowed to use their mobile phones, so their phones simply rang and vibrated during the interruptions.
@@
Each learner completed a single learning session on HTML from the WebApp MOOC. The session lasted an average of 30 minutes, with a maximum limit of 40 minutes. The tasks performed during the session were as follows:
@@
  • Calibration and Learner Information: Learners were briefed again on the session’s procedures. Sensors, such as the eye tracker and EEG band, were fitted to the learners and then calibrated. Learners were asked to provide information that could affect the sensors, such as the use of glasses/contact lenses, cardiac conditions, handedness (right or left-handed), etc. The specific medical conditions reported by the learners, including their prevalence, are summarized in Table 1.
  • EDA Test: Before starting the session, an initial EDA test was conducted to understand the learner’s stress levels. Learners were also asked to self-report their current state as "Very stressed, stressed, normal, relaxed, extremely relaxed" to obtain their feedback.
  • Form and Pretest: Additionally, they fill out a form with the following 2 questions, i) What is your current level of stress/anxiety? and ii) What is your current level of distraction? This aims to gather learner feedback before the session begins and to validate it with sensor data. We designed a pretest consisting of 8 multiple-choice questions of low to intermediate difficulty about HTML, with two objectives: i) to assess the learners’ current level of HTML knowledge and ii) to balance the three groups.
@@
The HTML unit consisted of a theoretical part where learners were introduced to HTML concepts in web design through explanatory videos and reading web documents, and a second part where learners’ understanding was evaluated. This evaluation involved 8 multiple-choice questions. After finishing, learners were able to review their incorrect answers. In neither part were learners allowed to return to a previous activity, pause the videos, or skip forward/backward in the video. The tasks were ordered as follows (see Fig. 2):
@@
    Videos: Two introductory videos on interactive web and basic HTML concepts. The first video lasted 20 seconds, and the second one lasted 1 minute and 48 seconds. During the second video, a message was sent to the learners in Groups 1 and 2 after 1 minute, coinciding with the explanation of HTML concepts.
@@
Table 2: Learner Feedback from Forms, Divided into Three Levels (Low, Medium, and High) and Expressed as Percentages.
@@
  • A final EDA Test was conducted to detect post-session stress levels after the session, and their feedback was collected again.
  • Learners completed a post-session Form, which asked about their stress, distraction, difficulty, and performance during the session. Additionally, learners in Groups 1 and 2 were asked about phone distraction, while Group 2 and 3 were asked about stress from phone deprivation. Table 2 presents the responses collected from the self-reported forms.
  • Learners were asked to fill out a form with the following 6 questions:
@@
    The final question varied depending on the group to which the learner belonged:
@@
    • For Group 2 and 3: "How stressed were you by not being able to use your mobile phone?"
@@ 20 Hz - 30 Hz MP4 files with codec H264
Table 3: Summary of Sensors and Data Collection from the IMPROVE Database.
@@
Two different platforms were used to monitor the learners: the edX and edBB platforms [1, 14]. The edX platform is an online MOOC platform used to conduct the learning sessions. It also allows the capture of session metadata from the learners, such as test answers, start and end times of tasks, and more. The edBB platform was used to monitor behavioral and biometric signals that have been shown to contain relevant information for understanding learner learning [1, 14, 21, 8, 19, 23, 24] (see Daza et al. [1] for a video demonstrationbbbhttps://www.youtube.com/watch?v=JbcL2N4YcDM). A multimodal acquisition framework was designed, incorporating 16 sources of information/sensors (see Fig. 1 and Table 3):
@@
  • Cameras: 2 Logitech C170 cameras (side and overhead) operating at 20 Hz with a resolution of $640\times 480$, and one front-facing RealSense camera were used. The RealSensecccwww.intelrealsense.com/wp-content/uploads/2020/06/Intel-RealSense-D400-Series-Datasheet-June-2020.pdf camera contains one RGB camera and two NIR cameras, with dimensions of 90 mm length $\times$ 25 mm depth $\times$ 25 mm height. The NIR cameras are monochrome and sensitive in both the visible spectrum and NIR, following the sensitivity curve of the CMOS sensors. The three cameras were configured to operate at 30 Hz and $1280\times 720$ resolution, and depth images are obtained by combining their three image channels. See Fig. 1 for more details on the position of each camera.
@@
  • Smartwatches: 2 different smartwatches, the Huawei Watch 2 and the FitBit Sense, were used to monitor heart rate [25, 26, 14], stress level (EDA sensor) [25, 27, 28], inertial sensors (gyroscope and accelerometer) [29]. It was decided to use 2 different smartwatch models, each worn on a different wrist, so that the heart rate monitored by the devices could be compared. The heart rate measurement was set to 1 Hz for the Huawei Watch 2 and 0.2 Hz for the FitBit Sense. The FitBit Sense has one sensors that are not available in the Huawei Watch 2: the EDA sensor. The Huawei Watch 2 was worn on the dominant hand of the learner to detect the movements produced by mouse usage.
@@
  • EEG: A NeuroSky EEG headsetdddhttps://store.neurosky.com/pages/mindwave, which measures the power spectrum density across 5 electroencephalographic channels ($\alpha,\beta,\gamma,\delta,\theta$). Through the preprocessing of these EEG channels, estimates of attention and meditation levels, as well as the occurrence of eye blinks, are obtained. EEG captures voltage signals typically generated by synaptic excitations of the dendrites in the pyramidal cells located in the top layer of the brain cortex [30]. These signals are primarily produced by the synchronous firing of numerous neurons and fibers [31]. EEG is an important, effective, and objective measure for estimating cognitive load and learner states [32, 33] because these signals are sensitive to the mental effort/cognitive works and mental states such as learning, lying, perception and stress. Extensive research in neurology and learning fields has explored the relationship between EEG signals and mental activities [23, 34, 31, 35, 36, 37], demonstrating the potential of the five electroencephalographic channels and the correlation between eyeblinks and cognitive load [38, 21, 23, 39].
@@
  • Visual attention: A Tobii Pro Fusioneeehttps://go.tobii.com/tobii-pro-fusion-product-description equipped with two high-speed infrared cameras configured at 120 Hz for eye tracking. This device estimated the following data: gaze origin and point, pupil diameter, eye movement type (fixation, saccade, unclassified, eyes not found), event duration, data quality, eyeblink, and more; allowing us to measure visual attention. Previous studies have demonstrated that visual attention is a key factor in learner learning, providing a deeper understanding of learner performance [40, 41]. There is a strong correlation between learner performance and visual attention, with features such as fixations, pupil diameter, eyeblink and saccades being closely linked to learning outcomes [40], fatigue detection [42], and task prediction [43].
@@
    Figure 3: Overview of the platforms and tools used to capture the IMPROVE database. The edBB setup was used to monitor the learner’s biometric and behavioral signals during the online session. The edX platform was utilized for delivering the course content and collecting session metadata. The LOGGE Tool was employed to record additional information, such as medical issues and group assignments, and monitor mobile phone usage. The data were aggregated into the IMPROVE database, which can be analyzed using the M2LADS platform to visualize and synchronize the captured signals and events.
@@
  • Computer Information: Relevant information for the session is obtained from various sources on the learner’s computer. i) Microphone Information: Audio is captured from the computer’s microphone at a frequency of 8000 Hz. The audio is recorded throughout the session, and all learners pronounce an initial phrase, "Sound test for the edBB platform," which can be used as a baseline for future research. ii) Keyboard and Mouse Activity: Keystrokes, mouse position, time between keystrokes, mouse wheel movements, etc., are monitored. Keystrokes and mouse dynamics have proven to be effective mechanisms for understanding learner sessions and providing secure certification [44, 45]. iii) Screen Capture: The monitor screen is captured at a frequency of 1 Hz. iv) Session metadata: Information such as keyboard type, logging data, IP and MAC addresses, operating system, etc., is collected.

### v3

# A multimodal dataset for understanding the impact of mobile phones on remote online virtual education

@@ Abstract
This work presents the IMPROVE dataset, a multimodal resource designed to evaluate the effects of mobile phone usage on learners during online education. It includes behavioral, biometric, physiological, and academic performance data collected from 120 learners divided into three groups with different levels of phone interaction, enabling the analysis of the impact of mobile phone usage and related phenomena such as nomophobia. A setup involving 16 synchronized sensors—including EEG, eye tracking, video cameras, smartwatches, and keystroke dynamics—was used to monitor learner activity during 30-minute sessions involving educational videos, document reading, and multiple-choice tests. Mobile phone usage events, including both controlled interventions and uncontrolled interactions, were labeled by supervisors and refined through a semi-supervised re-labeling process. Technical validation confirmed signal quality, and statistical analyses revealed biometric changes associated with phone usage. The dataset is publicly available for research through GitHub and Science Data Bank, with synchronized recordings from three platforms (edBB, edX, and LOGGE), provided in standard formats (.csv, .mp4, .wav, and .tsv), and accompanied by a detailed guide.
@@
Education is a fundamental pillar of society and has continually evolved throughout history. In recent years, it has been deeply influenced by the digital revolution, which has moved our society from the physical to the digital world. Evidence of this change in recent decades includes the growing significance of e-learning or online education, expected to grow exponentially over the next 20 years [1, 2]. Relevant institutions are adopting this model, especially those offering MOOCs (Massive Open Online Courses), due to their ability to reach a broader learner base [3].
@@
Learners have also been significantly impacted by the digital revolution, which has altered their habits to incorporate electronic devices into their daily routines such as mobile phones and tablets. Since 2012, studies [4, 5] have indicated that between 97% and 99% of learners own a mobile phone. Moreover, 95% of learners bring their phones to class every day, 92% use these devices to send text messages during class, and 10% even admit to texting during exams. On average, learners spend about five hours per day on their mobile phones, predominantly using messaging services (sending an average of 109 text messages per day) or social networks like WhatsApp and Instagram [6, 7].
@@
Therefore, the mobile phone raises numerous questions: Can it become a distraction to learners’ attention? Or, conversely, could the absence of a mobile phone cause greater distraction or anxiety? Considering that attention is defined as the ability to concentrate, specifically, to exert conscious cognitive effort on a particular task or stimulus at any given moment [8], and given that various studies suggest it is easy to lose focus due to short-term distractions such as messages, sounds, etc. [9, 10], it is reasonable to consider how mobile phones could impact learner performance. This is not a new idea, and various studies have shown that mobile phone use negatively affects learners’ attention, memory, and academic performance. This is mainly because it increases cognitive load by inducing learners to multitask [10, 5, 11, 12]. This effect becomes more noticeable as the class progresses or during long-duration activities [11]. However, learners’ performance can be influenced by their emotional state. For this reason, nomophobia, defined as the fear, stress, or anxiety caused by being unable to access or use a mobile phone, is an important factor to consider. Various studies have demonstrated how this condition affects memory and attention, ultimately impacting learners’ academic performance [13, 11].
@@
Interestingly, although nomophobia can lead to reduced academic performance, the previously mentioned studies have shown that just the presence of a mobile phone also diminishes learners’ attention levels, presenting a complex challenge that requires in-depth investigation.
@@
The IMPROVE dataset was collected in a realistic online learning environment. A total of 120 learners were monitored, divided into three distinct groups to understand the effects of mobile phone usage. To our knowledge, IMPROVE dataset is the first to provide a comprehensive multimodal acquisition setup including behavioral [1, 14], biometric, and learning analytics data from learners during an authentic online course that considers mobile phone usage. The dataset aims to facilitate understanding of the effects of mobile phone usage in online education. Additionally, it provides a robust dataset that can be utilized to develop new technologies to enhance learners’ learning processes or ensure secure certification [14, 15, 16, 17, 18, 19].
@@
We employed 16 sensors, recognized as effective indicators for understanding learner behavior [20, 8, 21, 22], including an electrodermal activity (EDA) sensor, an electroencephalogram (EEG) band, two pulse meters, an accelerometer, a gyroscope, three RGB cameras, two Near-Infrared (NIR) cameras, a mouse, a keyboard, a microphone, a monitor, and an eye tracker (see Fig. 1). These sensors enabled us to monitor cognitive activity, biometric, and behavioral signals throughout 120 online learning sessions, each lasting approximately 30 minutes. During these sessions, learners interacted with educational content related to HTML, including video lectures, document reading, and test completion to assess their knowledge both before and after the session (with a total of 16 questions per session). This comprehensive setup allowed us to track not only how learners engaged with the material but also how external factors, such as mobile phone interruptions, affected their focus and learning outcomes.
@@
In total, the dataset comprises 2.83 terabytes of data, making it one of the most extensive multimodal datasets in this domain. It includes 145 labeled mobile phone interruptions, providing crucial insights into how mobile phone usage affects learners.
@@
Figure 1: Acquisition setup used during data capture with the edBB[1, 14] platform, illustrating all sensors and devices employed, along with examples of the captured information.
@@ Methods
This section outlines the experimental methods and materials used for data collection. The dataset was obtained in accordance with the Declaration of Helsinki and was approved by the Ethics Committee of the Universidad Autónoma de Madrid (Approval No. CEI-130-2699, granted on 11/04/2023). Learners were informed about the experimental procedure and the data acquisition setup, and each participant signed a written consent form. The form clearly stated that the collected data would be used exclusively for research purposes and included in a publicly available dataset intended for the research community. Furthermore, participants gave explicit consent for the publication of non-anonymised videos showing their faces. Learners received an incentive for participating.
Dataset: IMPROVE
The dataset consists of 120 learners from the School of Engineering at the Universidad Autónoma de Madrid (UAM), who were monitored during a real online learning session focused on HTML, lasting between 20 and 35 minutes. The session was part of the MOOC titled “Introduction to Development of Web Applications” (WebApp MOOC for short) available on the edX platform (https://www.edx.org).
@@
Table 1: Distribution of learners by degree program, HTML proficiency levels, medical conditions (percentages are not shown because learners can have more than one condition), and gender, along with their average ages.
@@
Learners from the School of Engineering at the Universidad Autónoma de Madrid (UAM) were recruited. The School of Engineering was selected because its five degree programs all require basic knowledge of computer science, which aligns well with the online course. An economic incentive was offered to encourage participation. A total of 170 learners signed up for the study, of whom 120 were selected. An initial pilot test was conducted with six learners to verify the experimental protocol and monitoring system. Adjustments were made after the initial test, including: requiring Group 1 learners to provide more detailed responses to messages (for more information about the groups formed, see the subsection Protocol), giving learners a more comprehensive explanation of the session beforehand, removing an initial form to avoid potential suggestion effects, and making technical adjustments to the EEG band and the eye tracker calibration. The six learners from the pilot test did not participate in the final data collection. Four additional learners were excluded due to issues during the session and replaced by other participants. The final 120 learners were divided into three groups of 40. The selection process considered several criteria to ensure balanced groups, including gender (61 male, 59 female), HTML proficiency (none, beginner, intermediate, advanced), degree program, among other factors. Learners ranged in age from 18 to 32 years, with an average age of 21 (see Table 1). All learners gave their informed consent for the publication of the dataset. Each learner was anonymized with an identifier that cannot be linked to their identity.
@@
The objective was to examine the effects of mobile phone presence or absence on learner behavior and academic performance. To this end, three distinct groups were formed:
@@
    Group 1: Mobile phone use and possession were allowed. The device was placed on the learner’s desk, visible to the learner, with both sound and vibration activated.
@@
    Group 2: Mobile phone possession was allowed, while use was prohibited. The device was placed on the learner’s desk with the screen facing down, and sound and vibration were activated.
@@
    Group 3: The mobile phone was confiscated for the entire duration of the learning session.
Figure 2: Protocol followed during the learning session to evaluate the impact of mobile phone usage on learners’ academic performance. The top section presents the experimental design, illustrating the division of learners into three groups based on the permitted level of mobile phone interaction: unrestricted use, visible but prohibited use, and complete removal. The bottom section details the structure of the learning session, including the sequence of activities (e.g., video watching, reading, and completing assignments). It also shows a sample graph of a learner’s attention signal, synchronized with the session timeline. Time intervals in the signal are color-coded to reflect each activity, each accompanied by its corresponding icon, while windows of phone usage are highlighted by vertical boundary lines, with contextual icons indicating the type and timing of interruptions. This provides a visual representation of when phone-related interruptions occurred during the session.
To ensure consistency across all groups, all learners underwent the same learning session, delivered in a fixed sequence without enforced time limits between tasks. Learners were otherwise free to behave as they typically would in an e-learning context, with the only variation stemming from the mobile phone usage condition assigned to each group. No explicit instructions were given to avoid distractions unrelated to mobile phone use. However, two general restrictions applied: (i) learners could not return to previous tasks, and (ii) they were not allowed to control video playback (i.e., pause, rewind, or skip).
@@
Three supervisors were present throughout the session. The first supervisor was responsible for explaining the session in detail to the learner, describing the operation of the sensors, placing and calibrating them, and ensuring that the platforms functioned correctly. The second supervisor used the LOGGE tool [23] (for more details on this tool, see the "Data, platforms and sensors" subsection) to label the moments when the learner interacted with the mobile phone or when it rang, as well as to label other relevant information for the session. The third supervisor was responsible for sending two messages to the learners at key moments during the learning session for both Group 1 and Group 2, reviewing the information recorded by the second supervisor, noting any possible incidents during the session, and assisting the learner in case of doubts. We will differentiate between two different types of interruptions: interventions (the two controlled messages sent by the supervisor) and casual interactions (uncontrolled messages/calls received by the learner that were not related to the interventions). Therefore, interventions are considered controlled events, while interactions are classified as uncontrolled. The interventions were always delivered at the same points during the session. The first intervention occurred during an explanatory video of the lesson, and the second during the reading of HTML code, which was part of the learner evaluation (see Fig. 2). Learners in Group 1 were required to respond with a detailed reply. The messages consisted of simple questions such as “What do you think of the cafeteria of our School of Engineering?”, or an image from a popular movie like Harry Potter, accompanied by the question: “What do you think of this film series? Would you recommend it?” On average, learners took about 26 seconds to respond to the first message and about 36 seconds to respond to the second message. Learners in Group 2 were not allowed to use their mobile phones, so their phones simply rang and vibrated during the interruptions.
@@
Each learner completed a single learning session focused on HTML, taken from the WebApp MOOC. The session lasted an average of 30 minutes, with most learners completing it in 20 to 35 minutes, and a maximum duration of 40 minutes. The tasks performed during the session were as follows:
@@
  • Calibration and Learner Information: Learners were briefed again on the session procedures. Sensors, such as the eye tracker and EEG band, were fitted and calibrated. Learners were asked to provide information that could affect the sensors, such as the use of glasses/contact lenses, cardiac conditions, handedness (right or left-handed), etc. Specific medical conditions reported by the learners, including their prevalence, are summarized in Table 1.
  • EDA Test: Before starting the session, an initial EDA test was conducted to understand the learners’ stress levels. Learners were also asked to self-report their current state as "Very stressed, stressed, normal, relaxed, extremely relaxed" to obtain their feedback.
  • Form and Pretest: Additionally, they filled out a form with the following two questions, i) What is your current level of stress/anxiety? and ii) What is your current level of distraction? This information was collected to gather learner feedback prior to the session and to validate it against sensor data. Learners also completed a pretest consisting of eight multiple-choice questions of low to intermediate difficulty on HTML, with two objectives: i) to assess the learners’ current level of HTML knowledge and ii) to balance the three groups.
@@
The HTML unit consisted of a theoretical part where learners were introduced to HTML concepts in web design through explanatory videos and reading web documents, and a second part focused on assessing their understanding. This evaluation involved eight multiple-choice questions. After completing the test, learners were able to review their incorrect answers. At no point were learners allowed to return to previous activities, pause the videos, or skip forward/backward in the video. The tasks were ordered as follows (see Fig. 2):
@@
    Videos: Two introductory videos on interactive web and basic HTML concepts. The first video lasted 20 seconds, while the second lasted 1 minute and 48 seconds. During the second video, a message was sent to the learners in Groups 1 and 2 after 1 minute, coinciding with the explanation of HTML concepts.
@@
Table 2: Learner feedback from forms, categorized into three levels (low, medium, and high) and expressed as percentages. The responses cover aspects such as anxiety, distraction, perceived difficulty, and self-assessed performance, and are presented separately for each experimental group.
@@
  • A final EDA Test was conducted to detect post-session stress levels, and learners’ feedback was collected again.
  • Learners completed a post-session Form, which asked about their levels of stress, distraction, perceived difficulty, and performance during the session. Additionally, learners in Groups 1 and 2 were asked about phone distraction, while Groups 2 and 3 were asked about stress from phone deprivation. Table 2 presents the responses collected from the self-reported forms.
  • Learners were asked to fill out a form with the following six questions:
@@
    The final question varied depending on the group assigned to each learner:
@@
    • For Groups 2 and 3: "How stressed did you feel due to not being allowed to use your mobile phone?"
@@ 20 Hz - 30 Hz MP4 files with codec H264
Table 3: Summary of sensors and data collection from the IMPROVE dataset.
@@
Two different platforms were used to monitor the learners: the edX and edBB platforms [1, 14]. The edX platform is an online MOOC platform used to conduct the learning sessions. It also allows for the capture of session metadata from learners, such as test answers, task start and end times, and more. The edBB platform was used to monitor behavioral and biometric signals that have been shown to contain relevant information for understanding learner engagement and performance [1, 14, 22, 8, 20, 24, 25]. See Daza et al. [1] for a video demonstration (https://www.youtube.com/watch?v=JbcL2N4YcDM). A multimodal acquisition framework was designed, incorporating 16 sources of information/sensors (see Fig. 1 and Table 3):
@@
  • Cameras: Two Logitech C170 cameras (side and overhead) operating at 20 Hz with a resolution of $640\times 480$, and one front-facing Intel RealSense camera were used. The RealSense[26] camera includes one RGB camera and two NIR cameras, with dimensions of 90 mm length $\times$ 25 mm (depth) $\times$ 25 mm (height). The NIR cameras are monochrome and sensitive to both the visible spectrum and NIR, following the sensitivity curve of the CMOS sensors. The three Intel RealSense cameras were configured to operate at 30 Hz and a resolution of $1280\times 720$, and depth images were obtained by combining their three image channels. See Fig. 1 for more details on the position of each camera.
@@
  • Smartwatches: Two different smartwatches, the Huawei Watch 2 and the FitBit Sense, were used to monitor heart rate [27, 28, 14], stress levels via the electrodermal activity (EDA) sensor [27, 29, 30], and inertial data using gyroscopes and accelerometers [31]. Two different smartwatch models were used, each worn on a different wrist, so that the heart rate monitored by the devices could be compared. The heart rate measurement was set to 1 Hz for the Huawei Watch 2 and 0.2 Hz for the FitBit Sense. The FitBit Sense has one sensors that are not available in the Huawei Watch 2: the EDA sensor. The Huawei Watch 2 was worn on the dominant hand of the learner to detect the movements produced by mouse usage.
@@
  • EEG: A NeuroSky EEG headset, was used to measure the power spectral density across five electroencephalographic frequency bands ($\alpha,\beta,\gamma,\delta,\theta$). Through the preprocessing of these EEG channels, estimates of attention and meditation levels, as well as the occurrence of eyeblinks, are obtained. EEG captures voltage signals typically generated by synaptic excitations of the dendrites in the pyramidal cells located in the top layer of the brain cortex [32]. These signals are primarily produced by the synchronous firing of numerous neurons and fibers [33]. EEG is an important, effective, and objective measure for estimating cognitive load and learner states [34, 35] because these signals are sensitive to the mental effort/cognitive works and mental states such as learning, lying, perception and stress. Extensive research in neurology and learning fields has explored the relationship between EEG signals and mental activities [24, 36, 33, 37, 38, 39], demonstrating the potential of the five electroencephalographic channels and the correlation between eyeblinks and cognitive load [40, 22, 24, 41]. In the context of this dataset, EEG data were included to provide a direct and continuous measurement of learners’ cognitive states throughout the session. This information constitutes a key component for evaluating the impact of mobile phone usage on cognitive load, particularly with regard to fluctuations in attention levels and potential effects associated with nomophobia or mobile-induced distractions. This is especially relevant given that attention is defined as the cognitive effort directed toward a task and plays a pivotal role in ensuring accurate comprehension during learning.
@@
  • Visual attention: A Tobii Pro Fusion eye tracker equipped with two high-speed infrared cameras configured at 120 Hz for eye tracking. This device estimated the following data: gaze origin and point, pupil diameter, eye movement type (fixation, saccade, unclassified, eyes not found), event duration, data quality, eyeblink, and more; allowing us to measure visual attention. Previous studies have demonstrated that visual attention is a key factor in learning, providing a deeper understanding of learner performance [42, 43]. There is a strong correlation between learner performance and visual attention, with features such as fixations, pupil diameter, eyeblink and saccades being closely associated with learning outcomes [42], fatigue detection [44], and task prediction [45].
@@
    Figure 3: Overview of the platforms and tools used to capture the IMPROVE dataset. The edBB setup was used to monitor the learner’s biometric and behavioral signals during online sessions. The edX platform delivered the course content and collected session metadata. The LOGGE Tool recorded additional information, such as medical issues and group assignments, and mobile phone usage. The data were aggregated into the IMPROVE dataset, which could be analyzed using the M2LADS platform to visualize and synchronize the captured signals and events.
@@
  • Computer Information: Relevant information for the session were obtained from various sources on the learner’s computer. i) Microphone Information: Audio was recorded from the computer’s microphone at a sampling rate of 8000 Hz. The audio was recorded throughout the session, and all learners pronounce an initial phrase, "Sound test for the edBB platform," which can be used as a baseline for future research. ii) Keyboard and Mouse Activity: Keystrokes, mouse position, inter-keystroke timing, mouse wheel movements, etc., are monitored. Keystrokes and mouse dynamics have proven to be effective mechanisms for understanding learner sessions and providing secure certification [46, 47]. iii) Screen Capture: The monitor screen was recorded at a frequency of 1 Hz. iv) Session metadata: Information such as keyboard type, logging data, IP and MAC addresses, operating system, etc., was collected.